OpenStudy (anonymous):

Please help. The answer is -11/28, but I can't figure out how to get there. 2/3(x-7)=3/4(4x-5)

OpenStudy (tkhunny):

2/3(x-7)=3/4(4x-5) I'd be tempted to multiply by 12 8(x-7)=9(4x-5) Can you proceed from there?

OpenStudy (thomas5267):

Clever!

OpenStudy (anonymous):

I'm not sure .... sorry...

OpenStudy (thomas5267):

Do you know how to expand the brackets?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

I get 8x-42=36x-45

OpenStudy (thomas5267):

\(8\times7=56\) Now add multiples of x so that there is only x on one side and the constant on the other.

OpenStudy (thomas5267):

*add or subtract

OpenStudy (anonymous):

so it leaves me with 8x=36x-11

OpenStudy (anonymous):

then, x= 28x-11

OpenStudy (anonymous):

how do i get -11/28 out of that??